y(x + y)dx + (x + 2y − 1)dy = 0


1
Expert's answer
2021-04-25T07:24:11-0400

Solution.

"y(x+y)dx+(x+2y−1)dy=0,\newline
(yx+y^2)dx+(x+2y−1)dy=0.\newline
 \text{Let be } M(x,y)=yx+y^2, N(x,y)=x+2y-1."

Find integrating multiplier "\mu(x,y):"

"\frac{M_y-N_x}{N}=\frac{x
+2y-1}{x+2y-1}=1."

"\mu(x,y)=e^{\int 1dx}=e^x."

We will have

"(yx+y^2)e^xdx+(x+2y−1)e^xdy=0.\newline"

This equation is exact. Then "F(x,y)=\int (xy+y^2)e^xdx=y^2e^x+(xe^x-e^x)y+g(y),"

from here "F_y=2ye^x+xe^x-e^x+g'(x)=N(x,y)=(x+2y-1)e^x,"

then "g'(x)=0, g(x)=C," where "C" is some constant.

So, "e^xy^2+(xe^x-e^x)y+C=0" is the solution of equation.

Answer. "e^xy^2+(xe^x-e^x)y+C=0."